OpenStudy (anonymous):

Please help me!!!!! (-27x^3+54x)/9x

OpenStudy (lochana):

\[\frac{-27x}{9x} + \frac{54x}{9x}\]

OpenStudy (lochana):

can you do it now. something will cancel out

OpenStudy (anonymous):

Is that simplified?

OpenStudy (christos):

you missed x^3 in the numerator

OpenStudy (lochana):

ah yes. sorry about that

OpenStudy (lochana):

so \[\frac{-3*x^29x}{9x} - \frac{6*9x}{9x}\]

OpenStudy (lochana):

9x go away

OpenStudy (anonymous):

I'm confused...

OpenStudy (lochana):

and you get -3x^2-6

OpenStudy (lochana):

-27x^3 is -3.x^2.9x okay?

OpenStudy (lochana):

\[-27x^3 =-3.x^2.9x\]

OpenStudy (anonymous):

okay

OpenStudy (lochana):

and\[54x = 6.9x \]

OpenStudy (lochana):

so \[\frac{-27x^3 + 54x}{9x} = \frac{-3.x^2.9x + 6.9x}{9x} = -3x^2 + 6\]

OpenStudy (lochana):

and i did another mistake in my last answer. it should be +6. not -6

OpenStudy (lochana):

still confused?

OpenStudy (lochana):

call me back if you are still confused. type @lochana